### v2

Abstract
Ship trajectories from Automatic Identification System (AIS) messages are important in maritime safety, domain awareness, and algorithmic testing. Although the specifications for transmitting and receiving AIS messages are fixed, it is well known that technical inaccuracies and lacking seafarer compliance lead to severe data quality impairment. This paper proposes an adaptable, data-driven, $\alpha$-quantile-based framework for decoding, constructing, splitting, and assessing trajectories from raw AIS records to improve transparency in AIS data mining. Results indicate the proposed filtering algorithm robustly extracts clean, long, and uninterrupted trajectories for further processing. An open-source Python implementation of the framework is provided.
keywords:
[…]
1 Introduction and background
The emergence of the AIS represents a crucial advancement in maritime technology with implications for navigational safety and maritime domain awareness. The International Maritime Organization (IMO) played a crucial role in shaping the global adoption of AIS, incorporating mandatory requirements into the Safety of Life at Sea (SOLAS) regulation V/19. The required use of AIS systems for all ships of more than 300 gross tonnages on international voyages and cargo ships of more than 500 gross tonnages on national voyages became effective on 31 December 2004 (IMO, 2003).
Since then, a large group of researchers analyzed trajectories extracted from AIS records for various research domains such as anomaly detection (Pallotta et al., 2013; Zhang et al., 2016; Zhen et al., 2017; Rong et al., 2020; Wolsing et al., 2022), collision avoidance and risk assessment (Mou et al., 2010; Silveira et al., 2013; Chen et al., 2018; Rong et al., 2022; Zhang et al., 2023), pattern classification (Amigo Herrero et al., 2019; Sanchez Pedroche et al., 2020; Duan et al., 2022; Luo et al., 2023) or path planning (He et al., 2019; Xu et al., 2019; Gu et al., 2023; Waltz et al., 2023).
Concerning the methodology for trajectory extraction, a notable observation emerges. While a confluence of approaches exists, an agreed-upon standard within the scientific community is still absent. Trajectory extraction usually consists of two consecutive steps. First, filtering the data set on an individual message level to eliminate erroneous messages. Second, the cleaned messages are combined into meaningful, coherent trajectories.
Various procedures have been proposed to identify and eliminate abnormal individual AIS records. One filtering category applies to abnormal position reports, specifically those outside designated data areas. Zhang et al. (2018), Sang et al. (2015), Yuan et al. (2019), and Chen et al. (2020) have all explicitly addressed this issue. Another set of rules revolves around abnormal velocities. Let $SOG$ be a message’s speed over ground, then Yuan et al. (2019) suggests to exclude messages outside of $2kn<SOG<20kn$, while Sang et al. (2015) drops the current message if the absolute difference between the last and the current velocity is outside the interval $[0.3,0.8]$. Zhang et al. (2018) incorporates ship-type-specific maximum speeds, and Chen et al. (2020) sets a threshold at 30 knots for velocity-based exclusion. Abnormal Course over Ground (COG) is addressed by Yuan et al. (2019), Sang et al. (2015), and Chen et al. (2020), each proposing unique criteria, such as valid COG ranges and comparisons with theoretical tactical diameters. Abnormal Rate of Turn is considered by Zhang et al. (2018) and Yuan et al. (2019), with Zhang et al. (2018) specifying a maximum rotational limit based on vessel characteristics. Zhang et al. (2018) also introduces conditions for abnormal acceleration, restricting values within $[-1,1]$ knots per second, while abnormal jerk, the derivative of acceleration, is constrained by Zhang et al. (2018) to values beyond ±$\pm 15m/s^{3}$. Lastly, the issue of repeated data records is addressed by Yuan et al. (2019), albeit with unspecified criteria based on time judgments.
Following the initial filtration of individual messages, trajectories are typically constructed by sequentially linking messages emanating from the same source, identified by the Maritime Mobile Service Identity (MMSI) (Mao et al., 2018; Zhao et al., 2018; Yuan et al., 2019; Guo et al., 2021). Subsequently, an additional layer of trajectory-based filtering is applied to the generated trajectories. Several criteria have been proposed for filtering and splitting trajectories to enhance the accuracy and granularity of maritime data. Yuan et al. (2019), Zhao et al. (2018), and Mao et al. (2018) have contributed specific conditions for these purposes, while Yuan et al. (2019) suggests filtering trajectories based on the number of observations, proposing a threshold of 20 observations per track. Zhao et al. (2018) introduces a more stringent criterion, recommending a minimum of 100 observations for a track to be considered, whereas Mao et al. (2018) sets a threshold of 500 observations for tracks where the $SOG$ is not equal to zero, thereby emphasizing the importance of trajectories during active movement. Additionally, Mao et al. (2018) proposes a condition to avoid high-complexity trajectories, employing the average cosine of the angles between three consecutive message points as an indicator, with a threshold set at $0.8$.
Moreover, the concept of splitting trajectories into sub-tracks is addressed by Zhao et al. (2018). This involves three key conditions: a speed change greater than 15 knots between consecutive messages, a time interval exceeding 10 minutes between messages, and a permissible deviation between reported and generated time within $[-5,5]$ seconds, covering 96% of the ranges. These conditions collectively provide a comprehensive approach to delineating sub-tracks within longer trajectories, thereby extracting meaningful and discrete segments of vessel movement from AIS records.
A shared challenge in the aforementioned revised studies lies in determining diverse thresholds crucial for filtering records or trajectories. While certain thresholds are derived from empirical data, the majority are expert-derived, raising concerns about their generalizability and justification. In this paper, we advocate for an alternative approach to trajectory extraction, characterized by its reliance on data-driven methodologies with minimal reliance on predefined threshold values. Recognizing the inherent limitations and subjectivity associated with conventional threshold-setting, we aim to redefine trajectory extraction by minimizing assumptions and embracing a more empirically driven paradigm. Central to our approach is utilizing the $\alpha$-quantiles of empirical distributions constructed from different variables in the data, whose values will be used to distinguish coherent from incoherent trajectories.
Consequently, a comprehensive analysis was conducted on a dataset spanning 912 days of AIS records from January 2020 to June 2022. The geographical scope of this investigation encompasses the North Sea and regions of the Baltic Sea, with the primary objective of determining robust methodologies for extracting long, uninterrupted, and clean trajectories from the dataset. By clean, we understand the absence of erratic and highly volatile movement sequences of navigational parameters (SOG, COG, Position). All methodology presented in this article is publicly available as an open-source package (Paulig, 2024) for the Python programming language.
The rest of the paper is organized as follows. Section 2 briefly revisits AIS records and their structure, and Section 3 details the obtained data set. Section 4 introduces the split-point procedure based on empirical distribution functions, whose properties are detailed in Section 5. Section 6 proposes tools for assessing spatial features of the extracted trajectories, while Section 7 discusses the obtained results. Section 8 concludes. Additionally, C provides exemplary usage of the accompanying Python package.
2 AIS specifications
Over two decades ago, the ITU (2001) defined 27 different message types broadcast to serve many purposes, such as dynamic and static information about a voyage, safety-relevant information, or UTC inquiries.
AIS broadcasts are sent from two types of transceivers (Class A and Class B), dependent on ship size, purpose, and local regulations; where Class A transceivers meet the performance standards and carriage requirements adopted by IMO, while Class B only partially fulfill them. Class A stations autonomously transmit their position (message types 1/2/3) every 2-10 seconds, depending on the vessel’s speed and course changes, and at intervals of every three minutes or less when the vessel is anchored or moored. Additionally, these stations broadcast the vessel’s static and voyage-related information (message type 5) every 6 minutes.
Like Class A stations, Class B reports their position every three minutes or less when the vessel is anchored or moored. However, their position reports (messages 6 and 8) occur less frequently and with reduced power. Likewise, they transmit the vessel’s static data (messages 18/24) every 6 minutes, excluding voyage-related information. Due to the limited range and reliability, Class B reports are responsible for the small minority of records collected by base stations.
3 AIS data set
[…]
Figure 1: Geographical extent of the data set
(a) Count of messages by type.
(b) Absolute message count for each hour of the day, as calculated from receiving timestamps [UTC].
(c) Absolute message counts per month of the year, as calculated from receiving timestamps.
Figure 2: Descriptive statistics for the entire data set. Note that for the year 2022, the months from July until December are not part of the data set, and thus are not part of the absolute counts above.
The AIS data used in this study has been provided by the European Maritime Safety Agency (EMSA) and spans the geographical coordinates from $51.85^{\circ}$ to $60.49^{\circ}$ latitude and $4.85^{\circ}$ to $14.3^{\circ}$ longitude (refer to Figure 1). Due to its geographical extent, the data set contains messages received by base stations of the following countries: Belgium, Denmark, Estonia, Finland, France, Germany, Iceland, Ireland, Latvia, Lithuania, Netherlands, Norway, Poland, Russia, Sweden, and the United States military bases in Europe. The temporal scope encompasses the period from January 1, 2020, to June 30, 2022, which leads to approximately 3.2 billion (3 175 199 49931751994993\,175\,199\,4993 175 199 499) messages encoded as raw AIVDM / AIVDO sentences (refer to Table 1) (ITU, 2001), originating from 101 166 unique MMSIs. Dynamic messages, including types 1/2/3/18, constitute around 81.7% of the dataset, while the remaining messages pertain to static voyage reports of type 5, as presented in Figure 2(a).
Descriptively, intra-day message patterns demonstrate an area-specific and predictable “workday” trend, with a heightened message frequency observed during Coordinated Universal Time (UTC) daytime hours (see Figure 2(b)). However, monthly messages present an atypical surge in message count during July (see Figure 2(c)), whose origin remains unknown.
Message Type AIVDM sentence(s)
Type 1 !AIVDM,1,1,,,13‘dUP0P [^0] GqBjMw‘im0wvD0000,0*01
Type 5 !AIVDM,2,1,7,B,539g?6T00000@8i6221HU=E8LU>222222222220j1h5334@P04hTQCADR,0*15
!AIVDM,2,2,7,B,0EQC‘888888880,2*27
Table 1: Example of raw AIVDM sentences from the data set for a type 1 and a type 5 AIS message. Note, that type 1 messages consist of a single sentence, while type 5 messages are transmitted in two distinct sentences. A sentence always starts with the introducer "!AIVDM".
The data set was provided as one-day chunks in the .csv format, with messages of types 1/2/3/18 as one file and messages of type 5 in a separate file. The variables of the raw data set can be found in Table 2.
Variable name Description
timestamp UTC timestamp of message reception
message_id Message type
latitude Latitude of the message
longitude Longitude of the message
raw_message AIVDM sentence
MMSI Maritime Mobile Service Identifier
originator Country abbreviation of the receiving base station
Table 2: Variables of the raw data set.
The raw AIVDM / AVIDO sentences were decoded using the open-source pyais (Morien, 2023) Python package according to IMO regulations (IMO, 2010). In the decoding process of the AIS messages, an initial filtering stage was implemented to address the issue of duplicate data. This phenomenon arises when a vessel’s AIS report is within the coverage area of multiple base stations, leading to the same message being received from different locations. To mitigate this, any messages that were identical in content and received from more than one base station within a time frame shorter than the AIS system’s minimum sending frequency $f_{min}=2s$ were excluded.
The decoded AIS messages, however, cannot be used directly for trajectory extraction as they contain notable inconsistencies (Bailey, 2005; Harati-Mokhtari et al., 2007). Therefore, individual messages are analyzed to identify and eliminate those with erroneous positional and speed data. Subsequently, these messages are aggregated to construct trajectories. Given that a single trajectory may comprise multiple sub-trajectories, a data-driven, rule-based system is used to segment them.
3.1 Position reports
In line with methodologies outlined by Zhang et al. (2018), Sang et al. (2015), Yuan et al. (2019), and Chen et al. (2020), we systematically exclude messages with geospatial coordinates falling outside the defined spatial region of the data set. This process also conveniently addresses the issue of erroneous positional data. By setting logical boundaries for latitude ($\operatorname{lat}$) and longitude ($\operatorname{lon}$) – specifically, $51.85^{\circ}\leq\operatorname{lat}\leq 60.49^{\circ}$ and $4.85^{\circ}\leq\operatorname{lon}\leq 14.3^{\circ}$ – any data points falling outside these ranges are automatically discarded. Notably, the data set sourced from the EMSA had already undergone this preliminary filtering step; therefore, applying this spatial boundary criterion resulted in removing zero additional messages from the data set.
3.2 Velocity reports
The data feature an extensive range of reported SOGs, of which only a fraction is relevant for trajectory extraction. The analysis of the SOG data from 2021, as presented in Figure 3, indicates a notable number of observations registering a speed of approximately 102 knots. This observation aligns with established standards in the AIS protocol, in which a reported speed of 102.3 knots is conventionally used to signify that actual speed data is unavailable. This value does not represent an actual speed measurement but serves as a placeholder in the data set when speed information cannot be provided or is not applicable.
Figure 3 also reveals that a small fraction of messages ($\approx 0.3\%$) report velocities exceeding 30 knots. This observation is significant given the economic ramifications of high-speed maritime travel, such as elevated fuel consumption and accelerated attrition of materials. As also documented by Notteboom and Cariou (2009), commercial vessels rarely surpass the 30-knot threshold due to these constraints. Primarily, only specific categories of ships, such as high-speed crafts and certain military vessels, regularly operate beyond this speed limit.
In light of these insights, in concordance with Chen et al. (2020), this study opts to exclude all messages reporting speeds above 30 knots, thus keeping 99.7% of the data. This decision is grounded in the understanding that such high speeds are atypical for most commercial maritime traffic and, thus, may not represent typical operational behaviors and could potentially stem from errors in the ship’s transceiving equipment.
Figure 3: Histogram of speeds for all messages received in 2021 ($n\approx 1.07\times 10^{9}$). Note that the ordinate is log-transformed to improve visualization due to the large range of counts.
Figure 4: Trajectory of the vessel with MMSI 211XXXX90 laying at anchor in the harbor of Emden, Germany.
A significant proportion of the recorded messages ($>71\%$) indicated speeds below one knot. This prevalence of low-speed signals in the data set can be attributed to several factors, the most notable of which is the inclusion of non-moving, i.e., anchored or moored vessels. Slow-moving construction vessels, which are mandated to transmit AIS records regardless of their motion status, also contribute to these messages. For instance, Figure 4 illustrates an instance of a vessel at anchor with its AIS transceiver operational. While minimal, the motion of this vessel is still captured and transmitted as part of the AIS data. Such behavior, characterized by minimal or no actual navigational movement, should be considered for exclusion from data analyses focusing on vessel trajectories. For this reason, in this study, we remove messages whose SOG is less than one knot.
4 Trajectory construction
After the preliminary cleaning, the next step is to assemble the remaining AIS records into trajectories using a two-step approach. In the first step, we identify potential split-points within continuous vessel trajectories, defined as pairs of consecutive messages that appear to be more appropriately classified as belonging to separate trajectories. The separation criteria are based on empirical quantiles of several metrics detailed below. In the second step, the concept of absolute change of course of a trajectory is introduced to provide a flexible examination tool of the split trajectories.
A series of definitions are introduced below to facilitate comprehension and minimize confusion in the subsequent sections of this paper.
4.1 Definitions
Let $m$ represent a single AIS message. Define the set $\mathcal{T}_{k}=\{m\,|\,\operatorname{mmsi}(m)=k\}$ as the collection of all messages from a vessel with MMSI being equal to $k$, where $\operatorname{mmsi}(m)$ is a mapping of the AIS message to its corresponding MMSI. For this study, we focus on a vessel’s trajectory, i.e., ordered time-position tuples represented by the ordered set:
$\mathcal{T}_{k}^{>}=\{m_{1},m_{2},\dots,m_{p}\},\,\operatorname{unix}(m_{1})\leq\operatorname{unix}(m_{2})\leq\dots\leq\operatorname{unix}(m_{p}),\,\forall m\in\mathcal{T}_{k}.$ (1)
This set consists of the vessel’s transmitted messages, ordered by their UNIX timestamps in seconds, extracted from the messages via the $\operatorname{unix}(m)$ function. Furthermore, define $\operatorname{lat}(m)$, $\operatorname{lon}(m)$, and $\operatorname{sog}(m)$ functions returning latitude, longitude, and SOG, respectively, from a message $m$.
Some analyses require the computation of velocities obtained from spatial and temporal data. Given the Earth’s oblate spheroid geometry, distance estimation between geographical coordinates cannot be performed using Euclidean geometry. Predominantly, two methods are employed: the Haversine formula and the Vincenty (1975) algorithm, offering superior precision (Mahmoud and Akkari, 2016). However, for the objectives of this study, such a level of accuracy is unnecessary, predominantly due to the significantly larger error margins inherent in original GPS data (Jankowski et al., 2021). In addition, the Haversine formula’s computational efficiency is notably higher, making it a more pragmatic choice. It is defined as
$\operatorname{hav}(m_{1},m_{2})=2r\arcsin\left[\sqrt{\sin^{2}\left(\frac{\Delta_{\phi}}{2}\right)+\cos\{\operatorname{lat}(m_{1})\}\cos\{\operatorname{lat}(m_{2})\}\sin^{2}\left(\frac{\Delta_{\lambda}}{2}\right)}\right],$ (2)
with lateral and longitudinal differences $\Delta_{\phi}=\operatorname{lat}(m_{2})-\operatorname{lat}(m_{1})$, $\Delta_{\lambda}=\operatorname{lon}(m_{2})-\operatorname{lon}(m_{1})$, and $r=6371km$ being earth’s volumetric radius. The formulas’ output is a distance in meters, while latitude and longitude are provided in radians.
Furthermore, let
$\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}=\frac{\operatorname{sog}(m_{i})+\operatorname{sog}(m_{i+1})}{2},$ (3)
be the average reported speed over the ground for two consecutive messages $\operatorname{sog}(m_{i})$ and $\operatorname{sog}(m_{i+1})$, and
$\widehat{SOG}_{m_{i}}^{m_{i+1}}=\frac{\operatorname{hav}(m_{i},m_{i+1})}{\operatorname{unix}(m_{i+1})-\operatorname{unix}(m_{i})}\cdot\frac{1852}{3600},$ (4)
be the estimated SOG in knots calculated from the change in position over time. The scalar constant $1852/3600$ is the conversion factor that translates $ms^{-1}$ to $kn$.
4.2 Trajectory splitting
Consistent with the abovementioned research, we use the MMSI to assign AIS messages to their originating vessels and order these messages using the UTC timestamp to formulate a trajectory. Nevertheless, this methodology exhibits significant shortcomings without further processing, as it often results in trajectories with numerous undesirable characteristics, which will be outlined in this section. Accordingly, this study advances the trajectory segmentation methodology proposed by Zhao et al. (2018), applying it to ascertain the continuity of a trajectory by evaluating whether two successive AIS messages should be attributed to the same navigational path via a split-point approach. Therefore, $\alpha$-quantiles of various distributions, detailed in the next Sections, are computed to decide whether to split the trajectory at a particular point. The procedure is consolidated in Algorithm 1. Throughout the next subsections, it is assumed that if a trajectory is split multiple times such that any of the sub-trajectories only consists of one message, this sub-trajectory is discarded.
4.2.1 Temporal difference
The most natural metric to investigate for split points is the temporal difference between two successive messages of one trajectory $\operatorname{unix}(m_{i+1})-\operatorname{unix}(m_{i})$. Revisiting the AIS specifications, the defined transmission interval for dynamic AIS messages ranges from two to ten seconds for class A equipment ($2-30$ seconds for class B equipment) for vessels not at anchor and 180 seconds for vessels at anchor. Therefore, it is expected to find an accumulation of temporal gaps around these thresholds in the data, which could then be used to split trajectories. To verify the assumptions, a histogram of the temporal differences between two successive messages inside each trajectory for July 2021 is depicted in Figure 5. A brief discussion on why only a single month of data was used in the calculation is provided in Section 5. The distribution clearly shows that our initial assumptions are untenable, as temporal differences vary significantly in the interval from $[2,360]s$ with an exceptional peak around the $360s$ mark, whose origin remains undisclosed. The only message types in the AIS specification transmitted at $360s$ intervals are safety-relevant messages, none in the data set, and static voyage reports, which had not been considered.
The findings of this study suggest that it is not advisable to segment vessel trajectories based on the transmission intervals prescribed by AIS standards. Exemplary, we propose using the $(1-\alpha)\cdot 100=95$ percentile of the empirical distribution of temporal gaps between AIS messages to identify temporal split points. Following this logic, a split-point in this data set is given when the temporal interval between two successive messages exceeds $392s$ (see also Figure 5).
Figure 5: Histogram of the gap size for two consecutive dynamic messages in one trajectory. Gap sizes larger than 500s are truncated for better visualization. The histogram was built with speed-filtered data.
4.2.2 Velocity changes
Given the defined transmission interval for dynamic AIS messages, the variance in reported speeds between two adjacent records is anticipated to be reasonably small. This expectation is based on the premise that substantial acceleration is atypical for commercial vessels. Therefore, a threshold value must determine whether two consecutive messages belong to the same trajectory.
Figure 6: Histograms absolute SOG first differences (left), and turning rates for July 2021.
Following the procedure from the preceding paragraph, the absolute speed change distribution across all vessel trajectories for July 2021 was obtained. The results are illustrated on the right side of Figure 6, showing that only $\alpha=5\%$ of speed changes in individual trajectories are larger than $2.6kn$, resulting in a possible acceleration/deceleration range from $\frac{2.6\,kn}{392s}=0.007\,kn/s$ to $\frac{2.6\,kn}{2s}=1.3\,kn/s$, which are reasonable values, even for larger ships. This study identifies two messages with more than $2.6kn$ of absolute speed change as a split point, cutting it into two sub-tracks at this position.
4.2.3 Turning rates
It is generally not anticipated that significant deviations in a ship’s turning rate will occur during navigation, as rapid directional shifts not only increase material wear and tear but also contravene the IMO’s routing rationale advocating for safe and straightforward routing (IMO, 2019). Consequently, the right side of Figure 6 presents a histogram of tuning rates for two consecutive messages from one trajectory for July 2021, which will be used for threshold determination. Turning rates are calculated by dividing the difference of COG changes of two consecutive messages by their temporal difference. Adhering to the same $(1-\alpha)\cdot 100=95$ percentile criterion as applied to SOG changes, the eligible range of turning rates between two messages lies in the interval $[-0.48^{\circ}/s,0.38^{\circ}/s]$.
The relatively low turning rates may be attributed to several factors that underline the principles of safe and efficient maritime navigation. As mentioned, the IMO’s guidelines advocate for straightforward and safe routing. These guidelines are designed to minimize the risk of accidents, enhance the safety of the ship’s voyage, and protect the marine environment. Ships often use predetermined routes optimized for safety and efficiency, reducing the need for frequent or sharp turns. Further, many modern ships have advanced navigation systems like autopilots and GPS-based tracking. These systems are programmed to execute changes in course in a controlled and gradual manner, adhering closely to the calculated optimal route. Using such technology helps maintain the consistency of movements and avoid abrupt directional changes. Thus, the small range of turning rates reflects the combined impact of ship design, navigational technology, regulatory compliance, and environmental strategy aimed at ensuring a smooth, safe, and efficient voyage.
Figure 7: Example of a vessel’s trajectory that will be cut into several sub-trajectories by the heading change split-point threshold. The route exerts a ferry-like pattern where most of its trajectory is inconspicuous except for the turning points (around message numbers 50 and 100), where the ferry presumably performs a turn-around maneuver or is moored.
4.2.4 Reported against positional velocity
Figure 8: Erroneous positional outlier detected via the difference between the average reported SOG from the messages and the calculated speed from positional and temporal data. Latitude and longitude had been normalized for data protection
Another deficiency of trajectories that cannot be captured with the three metrics above are erroneous position reports that lead to significant outliers, as seen in Figure 8. In those cases, there is usually no anomaly detectable in terms of speed, course, or temporal differences, which is why this study follows Zhao et al. (2018) and uses the difference between $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}$ and $\widehat{SOG}_{m_{i}}^{m_{i+1}}$ to determine the outliers. Figure 9(a) provides context on the magnitude of differences across the admitted speed range. It can be observed that there are substantial deviations of different magnitudes in both directions. As with the other metrics, an empirical distribution of the differences is depicted in Figure 9(b) to establish a data-driven threshold. The $(1-\alpha)\cdot 100=95$ percentile for the present data falls within the left-skewed interval $[-8.96,6.65]\,kn$. The skewness of the distribution is expected, as $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}$ is bounded at $30kn$ due to the speed filtering employed in Section 3.2, while $\widehat{SOG}_{m_{i}}^{m_{i+1}}$ is unbounded.
(a) Scatter plot of the average observed SOG ($\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}$) and the calculated SOG from positional and temporal data ($\widehat{SOG}_{m_{i}}^{m_{i+1}})$ for all trajectories of the month of July 2021
(b) Histogram of the distribution of $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}-\widehat{SOG}_{m_{i}}^{m_{i+1}}$
Figure 9: Details on the divergence of $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}$ and $\widehat{SOG}_{m_{i}}^{m_{i+1}}$
4.2.5 Distance between messages
The final metric to be considered in our analysis is the distance between two successive AIS messages. The speed filtering process described in Section 3.2 involves removing trajectory segments where the SOG falls below one knot. This exclusion results in significant spatial gaps between successive messages within a trajectory and potentially leads to substantial temporal gaps if the vessel continues sailing. If the vessel increases its SOG to over one knot, its messages appear again in the data. In these cases, the discrepancy between the calculated speed (derived from positional and temporal data) and the average reported speed might be insufficiently large to be identified as split-points, based on the $(1-\alpha)\cdot 100=95$ percentile criterion.
To additionally capture these special cases, we also label two consecutive messages a split-point if their distance in miles, as calculated via the Haversine formula, is greater than the $(1-\alpha)\cdot 100=95$ percentile of the distance distribution obtained for the month of June 2021, which sets the boundary to $1.17nm$, with $nm$ standing for nautical miles.
4.3 Re-joining trajectories
As pointed out by Zhao et al. (2018), rejecting singular anomalous points may lead to an accidental separation of a trajectory that logically belongs together. The example in Figure 10 illustrates such a case. The original trajectory (Figure 10(a)) consists of a single anomalous point that should be removed from the data. Our split-point procedure, therefore, looks at consecutive overlapping pairs of messages and decides to split them based on the empirical distribution functions of the above-derived metrics. In the example, this leads to the trajectory being split into three separate trajectories, of which one only consists of a single message (Figure 10(b)). Since single-message-trajectories cannot sensibly count as such, they get removed automatically. In an attempt to rejoin the two remaining trajectories, the last message of the first trajectory and the first message of the second trajectory are again judged by the split-point procedure. If no split point is found, the two trajectories are rejoined into one (Figure 10(c)).
(a) Original trajectory.
(b) After split-point procedure applied.
(c) After re-joining.
Figure 10: Exemplary rejoin procedure. The split-point method splits the original, erroneous trajectory (a) into three separate ones, of which one only has a single message (b). Trajectories are judged again by the derived thresholds from the split-point method and possibly get rejoined (c)—illustration inspired by Zhang et al. (2018).
Data: Trajectory to split $\mathcal{T}^{>}$, Empirical quantile functions for change in speed $\hat{q}^{\Delta_{SOG}}$, tuning rate $\hat{q}^{{ROT}}$, time difference $\hat{q}^{\Delta_{t}}$, distance $\hat{q}^{dist}$, and reported against calculated speed $\hat{q}^{\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu-\widehat{SOG}}$, quantile threshold $\alpha$ (0.05 for this study).
 /* Set the thresholds */
$s\leftarrow\hat{q}^{\Delta_{SOG}}(1-\alpha)$
$r\leftarrow[\hat{q}^{{ROT}}(\alpha/2),\hat{q}^{{ROT}}(1-\alpha/2)]$
$t\leftarrow\hat{q}^{\Delta_{t}}(1-\alpha)$
$d\leftarrow\hat{q}^{dist}(1-\alpha)$
[…]
for $(m_{i},m_{i+1})\in\mathcal{T}^{>}$
       if $\operatorname{sog}(m_{i+1})-\operatorname{sog}(m_{i})>s$ (Section 4.2.2)
                 or
             $\frac{\operatorname{cog}(m_{i+1})-\operatorname{cog}(m_{i})}{\operatorname{unix}(m_{i+1})-\operatorname{unix}(m_{i})}\notin r$ (Section 4.2.3)
                 or
             $\operatorname{unix}(m_{i+1})-\operatorname{unix}(m_{i})>t$ (Section 4.2.1)
                 or
             $\operatorname{hav}(m_{i+1},m_{i})>d$ (Section 4.2.5)
                 or
             $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}-\widehat{SOG}_{m_{i}}^{m_{i+1}}\notin b$ (Section 4.2.4)
       then
             Split the trajectory between $m_{i}$ and $m_{i+1}$
       end if
end
Algorithm 1 Split-point procedure.
4.4 Applications of the split-point procedure to data
To demonstrate the effectiveness of the proposed split-point methodology, Figure 11 showcases a collection of trajectories derived from segments of the Little Belt and harbors of Aabenraa and Assens in Denmark, both with and without the application of the split-point procedure. In the absence of this procedure (as depicted in Figure 11(a)), numerous trajectories exhibit anomalously large distances between consecutive messages. This results in unrealistic jumps in the trajectory paths, leading to instances where the trajectories erroneously appear to traverse over land. This behavior can arise if multiple vessels transmit messages from a single MMSI, which is a common occurrence (Wu et al., 2017; Zhao et al., 2018; Yan et al., 2020), if trajectories are interrupted by the SOG-filter constraint or if vessels deliberately disable their AIS transceiver and re-enable it at a different location. In the filtered version, as depicted in Figure 11(b), the application of the split-point procedure effectively eliminates these anomalies, thereby accurately reconstructing the underlying structure of the waterway network. Quantitative information about the procedure can be found in Figure 11(c).
The split-point procedure also allows for significant insights into the navigational behaviors of vessels across different regions and types. This technique is particularly effective in generating ship-type dependent density maps, which illustrate the varying concentrations of vessel traffic along established routes and can also be used to discover new routes. This procedure allows key navigational routes and traffic flow dynamics specific to various ship types to be analyzed effectively. The methodology’s utility and effectiveness are demonstrated in Figures 14 and 15, which present detailed density maps of vessel traffic routes in the area of this research. Detailed information for practitioners on how the accompanying Python package implements Algorithm 1 and is utilized to generate these results is described in C.
(a) Trajectories without split-point filtering.
(b) Trajectories after applying the split-point procedure from Section 4.2.
Number of Raw Messages 10 420 4091042040910\,420\,40910 420 409 Percentage of Duplicates 3.11% Number of Trajectories 672 803672803672\,803672 803
Number of Split-Points 2 257 02322570232\,257\,0232 257 023 Number of Rejoined Tracks 47 4924749247\,49247 492 Average Trajectory length $2.66$ $nm$
(c) Quantitative information on the split-point procedure.
Figure 11: Effects of the split-point filter applied to an area around the Little Belt in Denmark. Only the time frame from 10/07/21 - 19/07/21 has been used to avoid clutter.
5 Time scale of empirical distributions
The split-point methodology outlined in Section 4.2 requires the computation of five distinct empirical distributions, using their quantiles to establish thresholds. In this process, we explicitly used trajectory data from one month of AIS records as the basis for each distribution. Figure 12 displays the quantile functions for all metrics, each calculated using different amounts of trajectory data (given as days and months). Visually, these distributions appear somewhat similar, leading to the assumption that the data follows identical distributions over various time horizons.
To substantiate this visual assumption, we compare the quantiles calculated from various time horizons against each other using the procedure outlined in Wilcox et al. (2014) using $10^{6}$ samples and $1000$ bootstrap replications each. We deliberately chose this setup to test how much information is pertained when using less data compared to one-year data. Contrary to initial expectations, our tests indicate that the distributions of different time horizons differ.
The computed quantile values in Table 4 only show minor differences in absolute terms, which leaves leeway to the choice of the amount of data used to calculate the quantiles. However, due to the huge sample sizes (over $10^{6}$ observations) that lead to very narrow confidence intervals, almost all quantiles are tested statistically different. Ultimately, we use one month of data throughout the study as a practical trade-off between accuracy and computational burden.
Nonetheless, since variations may emerge for distributions derived from shorter or longer time spans, additional adjustments by practitioners are potentially necessary.
Figure 12: Comparison of quantile functions for the split-points metrics using different amounts of data. Please note the different ordinate and abscissa scalings.
6 Spatial properties of split trajectories
We have formulated various trajectories, each comprising a reliable and consistent series of messages, notably free from anomalies. Nonetheless, we have no tool to assess the spatial properties of the extracted trajectories, especially concerning our original aim of extracting long, clean, and uninterrupted trajectories.
Therefore, we need a tool that measures these three quantities. Upon closer inspection, we find that all trajectories that had undergone the split-point procedure in Section 4.2 are already uninterrupted. Addressing the length and cleanliness of trajectories, we adopt a dual approach to control the length and the spatial properties of the trajectories. Consistent with methodologies employed in prior studies (Yuan et al., 2019; Zhao et al., 2018; Mao et al., 2018), the length of the trajectories is quantified using the count of messages per track, $n_{\mathrm{msg}}$. While effective in gauging trajectory length, this metric does not provide insights into the trajectory’s shape or spatial extent.
To bridge this gap, we investigate the convex hull each trajectory forms, or more precisely, its area. Since the calculation of a convex hull’s area relies on Euclidean geometry to apply, we first have to project the spheroid space of Earth to a Euclidean space, which is achieved using the Universal Transverse Mercator (UTM) projection (Krüger, 1912). While no isometric map from the sphere to the plane exists, the UTM projection still provides high accuracy if not used directly at the poles (Karney, 2011). Now, let $U=\{u_{1},u_{2},\dots,u_{n}\}$ be the set of UTM-projected positions of a trajectory. The convex hull $C(U)$ is the set of all convex combinations of positions from $U$. Formally, it is defined as
$C(U)=\left\{\sum_{i=1}^{n}\lambda_{i}u_{i}\bigm{|}\lambda_{i}\geq 0,\sum_{i=1}^{n}\lambda_{i}=1\right\},$ (5)
and in practice, is obtained using the quickhull algorithm by Barber et al. (1996). In two-dimensional space, the convex hull $C(U)$ forms a regular polygon whose area $A^{C}$ can be easily obtained.
From investigating the number of messages in and the convex hull area of a trajectory, it is not immediately clear how these metrics connect to its cleanliness. To measure this influence, this study adapts the idea of average complexity by Mao et al. (2018). In its original form, it is defined as
$\bar{c}(\mathcal{T}_{k}^{>})=|\mathcal{T}_{k}^{>}|^{-1}\sum_{m\in\mathcal{T}_{k}^{>}}\frac{\mathbf{p}^{\top}\mathbf{q}}{\|\mathbf{p}\|_{2}\|\mathbf{q}\|_{2}}$ (6)
with
$\mathbf{p}=\begin{bmatrix}\operatorname{lat}(m_{i})-\operatorname{lat}(m_{i-1})\\ \operatorname{lon}(m_{i})-\operatorname{lon}(m_{i-1})\end{bmatrix},\quad\mathbf{q}=\begin{bmatrix}\operatorname{lat}(m_{i+1})-\operatorname{lat}(m_{i})\\ \operatorname{lon}(m_{i+1})-\operatorname{lon}(m_{i})\end{bmatrix}$
which can be interpreted as the average cosine of the angle that the connecting lines of three consecutive messages enclose over the entire trajectory $\mathcal{T}_{k}^{>}$. The usage of this formula, however, has some drawbacks. In the original paper by Mao et al. (2018), only routes with an average complexity of more than 0.8 were included in the selection of accepted trajectories. We argue that the term complexity in this application is misleading as higher values of average complexity correspond to trajectories with smaller average course changes between messages. Additionally, the resulting values can be difficult to interpret in a maritime context as they merely describe a mathematical property of the examined trajectory.
Conveniently, a small adaption allows us to mitigate these shortcomings and construct an easily interpretable metric for assessing the cleanliness of a trajectory. We define the average absolute change of the course as
$\bar{\Delta}(\mathcal{T}^{>}_{k})=\pi^{-1}\cos^{-1}[\bar{c}(\mathcal{T}^{>}_{k})]\times 180\quad\text{for}|\mathcal{T}^{>}_{k}|>3$ (7)
which gives us a good intuition about the magnitude of directional changes in one trajectory in degrees. For a value of $\bar{c}(\mathcal{T}^{>}_{k})=0$, i.e., minimum complexity, the average absolute change of the course calculates to $180^{\circ}$ reflecting the U-turn behavior, while for $\bar{c}(\mathcal{T}^{>}_{k})=1$ the average absolute change of the course $\bar{\Delta}(\mathcal{T}^{>}_{k})=0^{\circ}$, encoding a perfectly straight trajectory.
Figure 13 presents a pixel map illustrating the impact of the number of messages ($n_{\mathrm{msg}}$) and the convex hull area ($A^{C}$) on $\bar{\Delta}(\mathcal{T}_{k}^{>})$ of trajectories in degrees. This analysis aggregates trajectories from 2021, both before (a) and following (b) the implementation of the split-point procedure, calculating $\bar{\Delta}(\mathcal{T}_{k}^{>})$ for each pixel. Each pixel represents a bin that averages $\bar{\Delta}(\mathcal{T}_{k}^{>})$ for trajectories within specific $n_{\mathrm{msg}}$ and $A^{C}$ ranges. The analysis is limited to $n_{\mathrm{msg}}\leq 100$ and $A^{C}\leq 5\times 10^{4}$ to maintain relevancy, as values beyond these thresholds yield negligible additional insights.
The results indicate that unprocessed trajectories exhibit no discernible pattern, suggesting that $n_{\mathrm{msg}}$ and $A^{C}$ are ineffective metrics for understanding unprocessed trajectory data. Conversely, applying the split-point procedure enhances the diversity of observed routes (fewer white spots) and demonstrates a fast trend towards higher smoothness values, particularly as $A^{C}$ increases.
(a) Average values of $\bar{\Delta}(\mathcal{T}^{>})$ for unfiltered trajectories.
(b) Average values of $\bar{\Delta}(\mathcal{T}^{>})$ for trajectories after split-point filtering.
Figure 13: $100\times 100$ pixel map of $\bar{\Delta}(\mathcal{T}^{>})$ in degrees as a function of the number of messages per track ($n_{\mathrm{msg}}$) and convex hull area ($A^{C}$). Each pixel resembles a bin, showing the average of $\bar{\Delta}(\mathcal{T}^{>})$ for all trajectories falling inside it. (a) considers unfiltered trajectories, (b) used the same data but with the split-point procedure applied. White pixels indicate the absence of data. The black line at $n_{\mathrm{msg}}=3$ is the minimum number of messages a trajectory needs to comprise to calculate $\bar{\Delta}(\mathcal{T}^{>})$. The convex hull area is restricted to $5\times 10^{4}$ for this visualization, as there are no changes in patterns above this threshold.
7 Discussion
This study provides a framework for trajectory extraction from raw, encoded AIVDM sentences. A multi-level filtering and splitting procedure has been proposed to construct long, clean, uninterrupted trajectories. Particular emphasis lies on the data-driven nature of the splitting process of this study. While other approaches (Sang et al., 2015; Zhao et al., 2018; Zhang et al., 2018; Yuan et al., 2019; Chen et al., 2020), too, propose functioning extraction processes, many threshold values used in those studies are still expert-derived, questioning their generalizability beyond the locally tested data. This study, instead, proposes to generate an empirical cumulative distribution function for the metric under consideration and use its 95th percentile bounds to determine the threshold values.
The methodology developed in this study demonstrates effective performance and aligns to establish a self-explanatory, data-driven approach for trajectory extraction. However, it is acknowledged that the decision to utilize the 95th percentile of the distributions as threshold bounds may be perceived as somewhat arbitrary. This choice, while not without its rationale, invites further research, as many derived thresholds have no specific meaning regarding established AIS transceiving regulations or the maritime domain in general.
8 Conclusion
Maritime AIS data is inherently noisy due to technical inaccuracies, seafarer inattention, or lacking compliance. However, trajectories extracted from AIS messages are valuable for maritime surveillance, domain awareness, and algorithmic testing, for which we propose a data-driven, open-source trajectory extraction and refinement framework. The extraction process is based on a split-point procedure that uses quantile boundaries of five empirically derived distribution functions to decide whether two consecutive messages belong to the same trajectory or should be split. We found that this procedure effectively eliminates anomalies in the data without resorting to expert-derived thresholds, making it universally adaptable to any AIS data set. Nevertheless, carefully scrutinizing the quantile boundary and threshold values is recommended, as they may vary significantly depending on data availability. For spatial understanding, we introduced the concept of average absolute change of course, $\bar{\Delta}(\cdot)$, which allows us to assess the presence or absence of erratic positional movement in a trajectory and recommended the number of messages per trajectory ($n_{\mathrm{msg}}$) as well as the convex hull area ($A^{C}$) as means to control $\bar{\Delta}(\cdot)$. The entire extraction and assessment pipeline, described by this article, is publicly available as an open-source Python package at Paulig (2024).
9 Acknowledgments
We thank the Center for Information Services and High-Performance Computing at TU Dresden for providing its facilities for high throughput calculations. We thank the European Maritime Safety Agency for providing the raw AIS records analyzed in this study. Further, we would also like to thank Martin Waltz for his valuable support throughout this project.
Appendix A $A^{C}_{v}$ for different ship types
Given that $ship\_type(MMSI)$ is a function mapping an MMSI to its corresponding ship type $v$, and $A^{C_{\mathcal{T}_{k}^{>}}}$ is the convex hull area for trajectory $\mathcal{T}_{k}^{>}$, the ship-type-average of their convex hull area can be found via
$\bar{A}^{C}_{v}=|\mathcal{V}|^{-1}\sum_{\mathcal{T}_{k}^{>}\in\mathcal{V}}A^{C_{\mathcal{T}_{k}^{>}}},$ (8)
where $\mathcal{V}=\{\mathcal{T}_{k}^{>}|ship\_type(k)=v\}$ is the set of all trajectories belonging to one ship type.
[…]
CARGO 174 100174100174\,100174 100
TANKER 173 443173443173\,443173 443
NOTAVAILABLE 141 221141221141\,221141 221
TUGTOW 126 509126509126\,509126 509
PASSENGER 115 021115021115\,021115 021
MILITARY 109 111109111109\,111109 111
OTHER 99 7719977199\,77199 771
PLEASURE 93 9029390293\,90293 902
SAILING 79 7307973079\,73079 730
WIG 78 3847838478\,38478 384
FISHING 60 6106061060\,61060 610
HSC 37 6183761837\,61837 618
Table 3: Average convex hull area for all trajectories with $n_{\mathrm{msg}}>50$ for all ship types for the year 2021. The ship types are specified according to ITU standards (ITU, 2001). For this analysis, only the base ship type had been distinguished; for example, any tanker (types 80 - 89) is included in the TANKER category. NOTAVAILABLE refers to the default if no ship type has been transmitted. OTHER are all ships not fitting into the above categories.
Table 3 collects these averages. In line with our intuition, cargo ships’ routes cover the largest average area of all ship types, as one of their primary goals is to ship goods effectively over long distances. Interestingly, Tug and tow boats cover the 3rd highest convex hull area. This finding might initially seem paradoxical given their role in assistive navigation in ports or terminals, typically associated with short and compressed trajectories. However, this observation can be rationalized by considering that our analysis only includes trajectories with more than 50 messages and a minimum speed of $1kn$. A visual inspection using heatmaps in Figures 14 and 15 revealed that Tug and Towboats travel longer routes than expected under these circumstances. On the other hand, High-Speed-Crafts (HSC) are mostly used as high-speed ferries for short passages in commercial contexts, which explains their relatively small area covered.
(a) Cargo
(b) Fishing
(c) High Speed Crafts
(d) Military
(e) Not available
(f) Other
Figure 14: Heatmap of route densities for different ship types for the year 2021. Plots generated from raw AIS records using the split-point method and $A^{C}>3\times 10^{4}m^{2}$ and $n_{\mathrm{msg}}>50$-refinement
(a) Passenger
(b) Pleasure
(c) Sailing
(d) Tanker
(e) Tug Tow
(f) Wing in Ground (WIG)
Figure 15: Continuation of Figure 14
Appendix B Quantile comparison results
Metric Quantiles used in Section 4.2
Upper and lower quantiles
$q=0.025$ $q=0.975$
$D=1$ $D=7$ $D=30$ $D=120$ $D=1$ $D=7$ $D=30$ $D=120$
Difference between reported
and calculated speed $[kn]$
-3.577 -3.149 -2.866 -4.078 4.041 4.080 4.053 4.334
Turning rates $[{}^{\circ}/s]$0.507 -0.466 -0.466 -0.491 0.421 0.376 0.372 0.400
Only upper quantiles
$q=0.95$
$D=1$ $D=7$ $D=30$ $D=120$
Time difference
between messages $[s]$
397.0 420.0 430.0 419.0
Speed changes $[m/s]$ 2.800 2.800 2.799 2.700
Distance between
consecutive messages $[nm]$
1.147 1.142 1.110 1.123
Table 4: Quantile values for the metrics used in Section 4.2. $q$ is the value of the quantile, $D$ is the number of days of data used to obtain the quantiles.
Appendix C Python package implementation details
All pre-processing, filtering, extracting, splitting, assessment and visualization procedures outlined in this article are available as the open-source Python package PyTSA (Python Trajectory Splitting and Assessment Agent, Paulig (2024)). This Section is intended to provide implementation details and practitioners guidelines on how to use PyTSA to reproduce the results presented in this research. For installation instructions, please refer to the package documentation.
Decoding raw AIS messages
If the data to be analyzed is present as raw AVIDM/AVIDO sentences, they must first be decoded. To do this with PyTSA, the input files must be provided as .csv files sorted such that one group of files only contains dynamic messages of types 1/2/3/18 and the other only messages of type 5 (static voyage reports). Both files must be named identically by the day for which they hold AIS messages, separated by underscores ("YYYY_MM_DD.csv").
Subsequently, decoding can be performed via
{minted}
[linenos,frame=lines,fontsize=]python from pytsa import decode
decode( source = "path/to/raw_dir", dest = "path/to/decoded_dir", njobs = 1 )
which will take all .csv files in the "source" folder, decode them, and store the decoded files in the .csv format in the "dest" folder. For more information on the "njobs" keyword and the structure of the data files, the user is deferred to the package documentation.
Trajectory extraction using the SearchAgent
PyTSA’s central object, the SearchAgent, provides a convenient tool for constructing, filtering, and splitting trajectories extracted from the messages decoded earlier. The SearchAgent class is instantiated using at least three components:
  1. 1.
    mgs12318file - A single file path or a list of file paths for dynamic messages
  2. 2.
    mgs5file - A single file path or a list of file paths for static messages
  3. 3.
    frame - A PyTSA BoundingBox object determining the spatial extent of the extraction process.
If, instead, it is desired for the data to undergo a pre-processing step, a pre-processing function that all messages must pass through before extraction can be defined. For example, to apply the speed constraint from Section 3.2, the instantiation can be performed as described in Code Listing C.1.
{minted}
[linenos,frame=lines,fontsize=]python import pandas as pd import pytsa
from pathlib import Path
# Column name for SOG in the decoded # csv data. _speed = …
# Filter out all vessels that are slower than 1 knot # and faster than 30 knots. # Any pre-processor is mandated to take in a single # pandas DataFrame and return a single pandas DataFrame def speed_filter(df: pd.DataFrame) -> pd.DataFrame: return df[(df[_speed] > 1) & (df[_speed] < 30)]
# Bounding Box for the area of study in this research frame = pytsa.BoundingBox( LATMIN = 52.2, # [°N] LATMAX = 56.9, # [°N] LONMIN = 6.3, # [°E] LONMAX = 9.5, # [°E] ) # This can be either a single file or a # list of files dynamic_data = Path("/path/to/dynamic.csv") static_data = Path("/path/to/static.csv")
search_agent = pytsa.SearchAgent( msg12318file = dynamic_data, msg5file = static_data frame = frame, preprocessor = speed_filter # Apply the speed filter using the "preprocessor" keyword )
List of Code Listings C.1 Basic instantiantion of the SearchAgent class with a custom pre-processing function.
After creating the instance, the extract_all() method can be used to construct and, if desired, also split the trajectories according to the split-point procedure from Section 4.2. The method returns a Python dictionary with every unique MMSI as keys and their corresponding TargetShip objects as values. These objects contain information about the ship type and length and all trajectories identified as belonging to this MMSI. The method has an integrated switch to turn the split-point procedure on or off. If turned off, every returned TargetShip only contains a single trajectory, consisting of all AIS messages sent from its MMSI sorted by time.
We look at the code used to produce Figure 11 to demonstrate this approach’s capabilities. We also utilize the visualization module shipped with PyTSA for this approach. {minted}[linenos,frame=lines,fontsize=,mathescape]python # # This listing continues Code Listing C.1 # # To perform only trajectory construction without # using the split-point procedure, we use the # ‘skip_tsplit=True‘ option. ships = search_agent.extract_all(njobs=16,skip_tsplit=True)
# Define the area around Aabenraa AABENRAA = pytsa.BoundingBox( LATMIN = 54.9, # [°N] LATMAX = 55.5, # [°N] LONMIN = 9.2, # [°E] LONMAX = 10.0, # [°E] )
# Plot trajectories on the map for the area # around AABENRAA plot_trajectories_on_map( ships = ships, extent = AABENRAA )
# With split-point procedure applied ships = search_agent.extract_all(njobs=4,skip_tsplit=False) plot_trajectories_on_map( ships = ships, extent = AABENRAA )
Assessing spatial properties of extracted trajectories
[…]
# Construct a Recipe object to assess every trajectory based on the rule functions # given to it. In this case, we use the partial class to coerce the rules to only take a single argument, # which is the trajectory given to it. assessment = Recipe( partial(too_few_obs, n = 50), partial(convex_hull_area, area = 3e5) )
# Instantiate the Inspector and split the trajectories into accepted and rejected ones. # In this example, the rejected trajectories are discarded. accepted, rejected = pytsa.Inspector(data = ships, recipe = assessment).inspect()
# Save a 500x500 px heatmap for every ship type types = [t for t in pytsa.ShipType] names = [t.name for t in types] expanded = [] for t in types: if isinstance(t.value, int): expanded.append([t.value]) else: expanded.append(list(t.value)) for i,t in enumerate(expanded): a = mmsi:s for mmsi,s in accepted.items() if any(st in t for st in s.ship_type) binned_heatmap(targets = a, bb = frame, npixels = 500, title = f"Heatmap for names[i]")